OpenStudy (anonymous):

show all work to simplify 2/x-2/x-1+2/x-2

OpenStudy (anonymous):

Use complete sentences to explain how to simplify this expression. Remember to list all restrictions.

OpenStudy (anonymous):

Help me?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Get common denominators

OpenStudy (anonymous):

Ok.

OpenStudy (luigi0210):

Wait, could you write out the equation using the editor?

hero (hero):

I'm going to take a wild guess... Is it: \[\frac{2}{x} - \frac{2}{x-1} + \frac{2}{x-2}\]

OpenStudy (anonymous):



hero (hero):

That's a bit different from what you wrote the first time @Littlemsadventuretime

hero (hero):

You seemed to have switched signs for the last two terms.

OpenStudy (anonymous):

sorry

hero (hero):

Okay, so which post is correct?

OpenStudy (anonymous):

the second

OpenStudy (anonymous):

is the correct eqaution

hero (hero):

An equation would have an equal sign.

OpenStudy (anonymous):

problem i mean

hero (hero):

To add or subtract fractions, the denominators must be the same.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

I just dont know